Parallel to x - y = 4 and passing through (-3,2) Express in standard form
Answer by algebrahouse.com(1659)

Parallel lines have equal slopes

Slope-intercept form is y = mx + b
m is the slope
b is the y-intercept

x - y = 4
-y = -x + 4 {subtracted x from each side}
y = x - 4 {divided each side by -1}
slope is 1

point is (-3,2) slope is 1
y - y1 = m(x - x1) {point-slope form}
y - 2 = 1[x - (-3)] {substituted into point-slope form}
y - 2 = x + 3 {simplified}
-x + y = 5 {put into standard form}

For more help from me, visit: www.algebrahouse.com